OpenStudy (anonymous):

If f(t)=t^3 +2, g(x)=2x+3, p(x)=(f(g(x)), and r(t)=f(g(t)). Compute r(-2).

OpenStudy (ranga):

r(t) = f(g(t)) r(-2) = f(g(-2)) g(x) = 2x + 3 g(-2) = ? Then put that number you find in the previous step in f(?) and evaluate it.

OpenStudy (anonymous):

g(-2)=-1 f(-1) = (-1)^3 +2 =1

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

so r(-2)=1?

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (ranga):

You are welcome.